what are the two answers for this problem x^3+512=0
Answer by oberobic(2304)

x^3 = -512
x = (-512)^(1/3) the cube root of -512.
.
x = -8
.
Always check your answer:
-8 * -8 = 64
-8 * 64 = -512
Correct.
.
Answer: -8 is the answer. There is no second answer.
.
Done.